In the United States, 200 million gallons of used motor oil is improperly disposed of each year. One gallon of used oil can contaminate one million gallons of drinking water. How many gallons of drinking water can 200 million gallons of oil contaminate? Express the answer in scientific notation.
Answer by Theo(13342)

200 million gallons improperly disposed of each year.
1 gallon contaminates 1 million gallons of drinking water.

200 million = 200 * 10^6
1 million = 1 * 10^6

for every gallon of oil, 1 million gallons of drinking water are contaminated.

that means that 200 * 10^6 * 1 * 10^6 gallons of drinking water get contaminated very year.

200 * 10^6 * 1 * 10^6 = 200 * 1 * 10^(6+6) = 200 * 10^12 gallons of drinking water get contaminated every year.

If you take 200 * 1 * 10^12 and divide it by 200 * 10^6, you get 1 * 10^(12-6) = 1 * 10^6 gallons of drinking water contaminated by each gallon of oil.

Since this is where we started from, the numbers look good and you answer would be:

200 * 10^12 gallons of drinking water get contaminated every year by 200 * 10^6 gallons of oil improperly disposed of each year assuming each gallon of oil contaminates 1 * 10^6 gallons of drinking water.